OpenStudy (anonymous):

solve by substitution; 2 = 2y -x 23 = 5y - 4x

OpenStudy (anonymous):

2-2y=x Substitute this in the second equation: 23=5y-4(2-2y) 23=5y-8+8y 23+8=13y 31=13y 31/13=y Substitute value of y in 1st equation: 2=2(31/13) -x 2+x = 62/13 x=62/13 -2 x=36/13.